.1. Consider the following initial-value problem. y = e(1+B)* In(1 + y²), 0<t<1 y (0) = a +1 a) ( 15p.) Determine the existenα=2 β=2
Answer #1

solution :- Given initial value problem y= e(+B)t In City), ostel 400)= 0+1 { Where D=2, B=2 a) We have to determine the exi(b) We have to use Eulers method with h=0.25 then find y(0.50) Here to 20 t, -0.25 te=0.50 Yo =3, - 9. - Yo+hf(to ,40) (to %

I HOPE THIS ANSWER WILL BE HELPFUL TO YOU

PLEASE LIKE THIS ANSWER

THANK YOU